THE ASSYMETRIC N-O STRETCH FUNDAMENTAL BAND OF NITROMETHANE: FTIR JET SPECTRA AND ASSIGNMENTS OF THE LOWEST 3 INTERNAL ROTOR STATES

DAVID S. PERRY, \underline{LOU DEGLIUMBERTO}, University of Akron, Akron, OH 44325; WENDY HEISERMAN, REBECCA COOK, HOWARD D. METTEE, Youngstown State University, Youngstown, OH 44555; ROBERT L. SAMS, Pacific Northwest National Laboratory, Richland, WA 99352; GEORG O. SORENSEN, University of Copenhagen, DK-2100, Copenhagen, Denmark.

The rotationally resolved spectrum of the b-type band near 1584 cm^{−1} of nitromethane has been recorded using the Bruker-120HR FTIR coupled to a continuous slit-jet expansion at PNNL. The rotational temperature of the spectrum is approximately 15K with a spectral resolution of 0.0025 cm^{−1}. Assignments have been made for 649 transitions using ground state combination differences. For the m′ = 0 internal rotor state, 370 transitions reaching J′ ≤ 18 and K′_{a} ≤ 5 have been assigned. For the m′ = 1 state, 232 transitions with J′ ≤ 15 and K′_{a} ≤ 4 have been assigned. For the m′ = 2 state, 47 transitions with J′ ≤ 12 and K′_{a} = 1 have been assigned. The J′-levels in the K′_{a} = 5 in the m′ = 0 state are split by a perturbation.